OpenStudy (anonymous):

simplify: (x^2y^3z^6)^0

OpenStudy (callisto):

a^0 = 1

OpenStudy (callisto):

Hmm... Got it?

OpenStudy (anonymous):

no :( i hate this kind of stuff

OpenStudy (callisto):

That's a property of exponential function :|

OpenStudy (anonymous):

would you multiply all of the exponents by 1 and it would just be the same?

OpenStudy (callisto):

 Exponential function has this graph.. They (all) pass through (1,0)

OpenStudy (anonymous):

@sauravshakya helpppppp lol

OpenStudy (callisto):

\[(x^2y^3z^6)^0 = (x^{2\times 0})(y^{3\times 0})(z^{6\times 0}) = x^0y^0z^0=...\]

OpenStudy (anonymous):

so it would just be xyz?

OpenStudy (callisto):

a^0 = 1 ... :| So, x^0 =1... Guess what y^0 and z^0 are!

OpenStudy (anonymous):

1? lol would you multiply all those to still = 1 or would you add them all?

OpenStudy (callisto):

1... \[x^0y^0z^0 = 1\times 1 \times 1\]

OpenStudy (anonymous):

ok thats what i thought

OpenStudy (callisto):

It's true when x, y, z >0

OpenStudy (anonymous):

uhg this is so confusing.. i would know this stuff if i was at school.. but i was sick the day we did this stuff

OpenStudy (callisto):

An easier way is to treat x^2y^3z^6 = a And apply the properties a^0 =1 to get the answer immediately ... Oh.. I'm sorry to hear that :(

OpenStudy (anonymous):

so the whole thing would = 1?

OpenStudy (callisto):

*property Yes.

OpenStudy (callisto):

Same answer as you get from \((x^2y^3z^6)^0 = (x^{2\times 0})(y^{3\times 0})(z^{6\times 0}) = x^0y^0z^0=...\)

OpenStudy (anonymous):

so i could say x,y,z = 1? or just 1

OpenStudy (callisto):

No! It's \((x^2y^3z^6)^0 =1\) That doesn't mean x, y, z =1 :\